Consider the following balanced chemical equation:

2H2S(g) + SO2(g) → 3S(s) + 2H2O(g)

a. How many moles of H2S must react to form 4.18 moles of S?

_ moles H2S

b. How many moles of SO2 must react to form 4.18 moles of S?

_ moles SO2

c. How many moles of H2S must react to form 4.18 moles of H2O?

_ moles H2S

d. How many moles of SO2 must react to form 4.18 moles of H2O?

_ moles SO2
Answer #1

a)

From reaction,

Mol of H2S reacting = (2/3)*moles of S

= (2/3)*4.18 moles

= 2.79 moles

Answer: 2.79 moles

b)

From reaction,

Mol of SO2 reacting = (2/3)*moles of S

= (1/3)*4.18 moles

= 1.39 moles

Answer: 1.39 moles

C)

From reaction,

Mol of H2S reacting = (2/2)*moles of H2O

= (2/2)*4.18 moles

= 4.18 moles

Answer: 4.18 moles

D)

From reaction,

Mol of SO2 reacting = (1/2)*moles of H2O

= (2/2)*4.18 moles

= 2.09 moles

Answer: 2.09 moles